# Weiwei Ding1,2, Chuanguang Yang1,2, Yongjun Xu1,2, Xiaolong Jin1,2,3, Jiafeng Guo1,2,3 1

### v1

@@ Abstract
Knowledge Base Question Answering (KBQA) aims to answer natural-language questions over a structured Knowledge Base (KB). Recent work improves KBQA by adopting an agentic reasoning paradigm, in which Large Language Models (LLMs) iteratively decompose a question, generate its corresponding logical queries, and interact with the KB to derive the answer. However, these methods typically fine-tune LLMs on reasoning trajectories synthesized via process supervision, which offers weak incentives for exploration and thus fails to strengthen the agentic reasoning ability. In this paper, we propose KnowCoder-A1, an LLM that can autonomously perform agentic reasoning on KBs to obtain answers. To incentivize autonomous exploration, KnowCoder-A1 trains the LLM under outcome-only supervision via a multi-stage curriculum reinforcement learning with an easy-to-hard curriculum. To establish foundational agentic capabilities, KnowCoder-A1 first fine-tunes the LLM on a small set of high-quality trajectories obtained through outcome-based rejection sampling. Then, to alleviate the reward sparsity inherent in outcome-only supervision, it applies multi-stage curriculum RL with reward schedules that progress from easy to hard. Trained with outcome-only supervision, KnowCoder-A1 exhibits powerful reasoning behaviors and consistently outperforms prior approaches across three mainstream datasets. Notably, on the zero-shot subset of GrailQA, KnowCoder-A1 achieves up to an 11.1% relative improvement while using only one-twelfth of the training data, demonstrating strong agentic reasoning capabilities.
@@ 1 Introduction
Knowledge Base Question Answering (KBQA) leverages the rich semantic information in Knowledge Bases (KB) to deeply understand natural language questions and provide precise answers. Since KBQA holds significant value in applications such as search engines, intelligent healthcare, and financial risk control (Jang et al., 2017; Wu et al., 2024; Chen et al., 2025), it has garnered extensive attention in recent years. Although KBQA performs well on simple questions, it still struggles with complex questions in real-world scenarios. This difficulty primarily arises from generalizing to a wide range of unseen complex questions and adapting to the heterogeneous structures of KBs.
@@
Existing semantic-parsing-based (Shu et al., 2022; Jiang et al., ) and information-retrieval-based KBQA approaches (Liu et al., 2022; Sun et al., 2019) rely on predefined workflows (e.g., generate-then-retrieve or retrieve-then-rank pipelines), resulting in error propagation and adaptability limitations. Recently, a series of approaches has adopted an agentic reasoning paradigm. As illustrated in the middle of Fig. 1, under this paradigm, LLMs autonomously perform task planning, tool calling, and ultimately derive precise answers through multi-turn interactions with KBs. Specifically, existing agentic approaches fall into two categories, i.e., prompting-based (Xiong et al., 2024; Sun et al., ) and finetuning-based approaches (Jiang et al., 2024; Fang et al., 2024; Zhu et al., 2025). The former uses hand-crafted few-shot prompts to guide reasoning, relying on the intrinsic reasoning ability of LLMs. However, few examples are inadequate for the model to understand complex questions, perceiving the KB structure, resulting in low performance. In response, finetuning-based approaches adapt open-source LLMs by constructing agentic reasoning trajectories and fine-tuning the model to follow them. However, most of them rely on process supervision. Since manually creating high-quality trajectories is labor-intensive, they typically synthesize reasoning trajectories by decomposing gold logical forms (e.g., SPARQL). Besides, they learn from trajectories via Supervised Fine-Tuning (SFT) to employ the model with agentic reasoning ability.
@@ 2.1 Traditional KBQA methods
Traditional KBQA methods can be categorized into two primary kinds of approaches: Information Retrieval-based (IR-based) and Semantic Parsing-based (SP-based) ones. IR-based methods (Miller et al., 2016; Liu et al., 2022; Sun et al., 2019; Saxena et al., 2020; Shi et al., 2021; Zhang et al., 2022) first retrieve query-relevant information from the KB and subsequently rank candidate answers. Among IR-based methods, some (Bordes et al., 2015; Chen et al., 2019a) incorporate a long-term memory component that can be read from and written to, or perform a path-walking process over the KB to derive the answer (Qiu et al., 2020; Ren et al., 2021). SP-based methods (Sun et al., 2020; Chen et al., 2019b; LAN & JIANG, ; Jiang et al., ; Ye et al., 2022; Shu et al., 2022) transform the natural language question into a logical query, link the entities and relations, and execute the completed query to obtain the answer. With the advent of LLMs, they have been employed to enhance both approaches, serving as powerful subgraph retrievers (Zhao et al., 2024; Liu et al., 2024) or semantic parsers (Li et al., 2023; Nie et al., 2024). However, these enhancements do not fundamentally alter the predefined, multi-stage workflow, which brings inflexibility and error propagation.
@@ 2.2 Agentic KBQA methods
Recently, a new paradigm of agentic reasoning has emerged, which views the LLM as an agent that uses tools to interact with the KB, progressively reason, and finally find the answer. We call them agentic methods in this paper. They can be divided into two main categories: prompt-based and finetuning-based ones. The former methods, such as Interactive-KBQA (Xiong et al., 2024), QueryAgent (Huang et al., 2024), Readi (Cheng et al., 2024), and TOG (Sun et al., ), leverage carefully designed interaction strategies and prompts to elicit the model’s reasoning capabilities. However, these methods struggle to equip the model with sufficient agentic capability to effectively perceive the KB and execute complex reasoning. Their success is also heavily dependent on the intrinsic capabilities of LLMs, most of which are closed-source and have an extremely large number of parameters. However, the limited number of examples is insufficient for the model to fully comprehend complex questions and capture the underlying KB structure, leading to suboptimal performance. Finetuning-based methods aim to distill reasoning abilities into smaller, open-source LLMs. Approaches (Gu et al., 2023; Jiang et al., 2024; Zhu et al., 2025; Fang et al., 2024) achieve this by constructing high-quality reasoning trajectories and finetuning on them. However, these trajectories, by decomposing and augmenting gold program annotations and the reliance on process supervision via SFT, limit the model’s capacity for autonomous exploration, resulting in poor robustness and flexibility. Although some recent works have used Monte Carlo Tree Search (MCTS) methods (Luo et al., ; Xiong et al., 2025) to expand samples during inference for incremental fine-tuning, this is an indirect way to foster exploration and still operates within the confines of the process-supervised SFT paradigm. Overall, existing finetuning-based methods fall under SFT-based process supervision, which, as noted in Section 1, struggles to encourage autonomous exploration, leading to less robust and adaptable reasoning trajectories.
@@ 3 Preliminaries
KBQA. Given a natural-language question $q$, the KB $\mathcal{G}$, and a set of topic entities $E_{q}\subseteq\mathcal{E}$ mentioned in $q$, the objective of the KBQA task is to find the answer set that maximizes the conditional probability $P(\mathcal{A}_{q}|q,E_{q},\mathcal{G})$. Notably, following prior work Luo et al. , we assume entity mentions in $q$ are linked to the knowledge graph $\mathcal{G}$ and given as an input.
@@ 4.1 Agentic Initialization
KnowCoder-A1 follows a ReAct-style (Yao et al., 2023) agent paradigm that interleaves explicit reasoning and tool use over a KB environment, and terminates with a final answer.
@@
Action Space. At each step, the agent first produces an explicit thinking process and then makes a tool call from a finite action set Action. We instantiate three tools based on Interactive-KBQA (Xiong et al., 2024):
@@ 4.2 SFT-based Cold-start Stage
Next, to endow the model with stronger exploratory abilities and expand its exploration space, we use a prompt designed to encourage trial-and-error. As the yellow block in Fig. 2 shows, the prompt contains three parts: task description, tool definition, and high-quality demos: a few manually crafted exemplars that demonstrate a general agentic reasoning process, including planning, interacting, and self-correction after errors. Then, we prompt strong models (e.g., Gemini 2.5 Flash  (Comanici et al., 2025) and DeepSeek-V3  (Guo et al., 2025)) to generate multiple candidate trajectories for each question.
@@
  • Exact-match Correctness. The Exact Match (EM) score (Talmor & Berant, 2018) between the predicted and gold answers must be $1$. This ensures that at least one of the predicted answers appears in the gold answers, which largely guarantees the correctness of the trajectory.
@@ 4.3 RL-based Exploration Stage
Based on the cold-start model, in the RL stage, we aim to further promote autonomous exploration and eliminate the reliance on process supervision. To achieve this, we employ the outcome-based Group Relative Policy Optimization (GRPO) method (Shao et al., 2024), where any trajectory leading to a correct answer receives a positive incentive, thereby encouraging the agent to explore and acquire stronger reasoning capabilities. Moreover, to ensure the RL training is both balanced and effective, we implement two key strategies.
@@ 4.3.1 Policy Optimization via GRPO
Inspired by the recent success of policy gradient methods on reasoning tasks (Jin et al., 2025; Xue et al., 2025; Luo et al., 2025), we employ the GRPO method for optimization. GRPO estimates the baseline using the rewards from a group of rollouts for the same query to compute the advantage. In this paper, we adopt a fully on-policy variant of GRPO, which is demonstrated as an effective way to incentivize exploration (He et al., 2025). Its objective function can be expressed as follows:
@@ 5.1 Experiment Setup
Dataset and Metrics. We evaluate KnowCoder-A1 on three widely-used KBQA datasets: WebQSP, CWQ, and the generalization-focused GrailQA, using a subset of the official testing data. Following existing work (Xiong et al., 2024), we primarily evaluate the F1 metrics, supplemented by Random Hits@1(RHits@1) and EM score. As our approach does not rely on semantic parsing, we do not report logical form accuracy (Gu et al., 2021).
Baselines. We compare KnowCoder-A1 against two main categories of methods: prompting-based and fine-tuning-based methods. For the former, we include a diverse range of prompting strategies, from direct prompting, such as IO, CoT, etc., and agent-based reasoning method Interactive-KBQA (Xiong et al., 2024). To ensure a fair comparison and reproducibility, we report the scores for these methods directly from their original publications. For the fine-tuning-based methods, we select SOTA methods finetuned on full data, include RnG-KBQA (Ye et al., 2022) and TIARA (Shu et al., 2022), and the directly comparable low-source methods, which are finetuned on low-source data, includes agentic approaches like SFT-traj. (Xiong et al., 2024), KBQA-o1 (Luo et al., ), and MCTS-KBQA (Xiong et al., 2025).
@@ A.1 Datasets
We evaluate the reasoning performance on three widely-used datasets: WebQSP (Talmor & Berant, 2018), CWQ (Talmor & Berant, 2018), and GrailQA (Gu et al., 2021). These datasets are constructed based on Freebase and consist of natural language questions paired with their corresponding SPARQL queries. Specifically, WebQSP features questions where the topic entity and answer entities are at most 2 relation hops apart. CWQ extends WebQSP by incorporating more complex entities and constraints, with relation paths up to 4 hops. In contrast to WebQSP and CWQ, GrailQA is designed for a more comprehensive evaluation of generalization capabilities across three settings: independently and identically distributed (I.I.D.), compositional, and zero-shot.
To ensure consistent results and reduce evaluation overhead, we follow the setting of previous work Xiong et al. (2024), which uniformly sample questions across different types in WebQSP and CWQ. For the GrailQA dataset, we randomly sample 600 instances for evaluation.
@@ A.2 Baselines
Fine-tuning-based Methods. We consider two settings: fine-tuning on partial data (low-source setting) and on full data. In the low-source setting, we select the agentic approaches, includes the SFT model in Interactive-KBQA (Xiong et al., 2024), which uses the same toolset as our method (abbreviated as SFT-traj.), two MCTS-based agentic methods, KBQA-o1 (Luo et al., ) and MCTS-KBQA (Xiong et al., 2025). Notably, since our method is trained on only a subset of the data and uses no SPARQL annotations, its performance is directly comparable to this category of baselines. For the full-data setting, which serves as a secondary point of reference, we include the SOTA SP-based methods RnG-KBQA (Ye et al., 2022) and TIARA (Shu et al., 2022).
Prompting-based Methods. We selected a diverse set of prompting strategies for comparison. These include methods based on direct prompting, Chain-of-Thought (CoT), and self-consistency with CoT; approaches that use an LLM to enhance semantic parsing (KB-BINDER(Li et al., 2023) and KB-Coder (Nie et al., 2024)) and subgraph retrieval (Tian et al., 2024); and Interactive-KBQA (Xiong et al., 2024) that prompts an LLM to reason as an agent. For a fair comparison, we report the scores for these methods directly from their official papers, as the specific API versions used in their experiments are no longer accessible.
@@ A.3 Evaluation Metrics
We primarily report the F1 score since the answers are returned as an unordered list. Following previous work (Xiong et al., 2024), we also report Random Hits@1 (RHits@1) and Exact Match (EM) for reference. As our method is not fully based on semantic parsing, we do not report the semantically equivalent of logical forms (another score commonly abbreviated as EM) in  Gu et al. (2021).

### v2

@@ Abstract
Knowledge Base Question Answering (KBQA) aims to answer natural-language questions over a structured Knowledge Base (KB). Recent work improves KBQA by adopting an agentic reasoning paradigm, in which Large Language Models (LLMs) iteratively decompose a question, generate its corresponding logical queries, and interact with the KB to derive the answer. However, these methods typically fine-tune LLMs on reasoning trajectories synthesized via process supervision, which offers weak incentives for exploration and thus fails to strengthen the agentic reasoning ability. In this paper, we propose KnowCoder-A1, an LLM that can autonomously perform agentic reasoning on KBs to obtain answers. To incentivize autonomous exploration, KnowCoder-A1 trains the LLM under outcome-only supervision via a multi-stage curriculum reinforcement learning with an easy-to-hard curriculum. To establish foundational agentic capabilities, KnowCoder-A1 first fine-tunes the LLM on a small set of high-quality trajectories obtained through outcome-based rejection sampling. Then, to alleviate the reward sparsity inherent in outcome-only supervision, it applies multi-stage curriculum RL with reward schedules that progress from easy to hard. Trained with outcome-only supervision, KnowCoder-A1 exhibits powerful reasoning behaviors and consistently outperforms prior approaches across three mainstream datasets. Notably, on the zero-shot subset of GrailQA, KnowCoder-A1 achieves up to an 11.1% relative improvement while using only one-twelfth of the training data, demonstrating strong agentic reasoning capabilities. Code is available at https://github.com/ICT-GoKnow/KnowCoder-A1.
@@ 1 Introduction
Knowledge Base Question Answering (KBQA) leverages the rich semantic information in Knowledge Bases (KB) to deeply understand natural language questions and provide precise answers. Since KBQA holds significant value in applications such as search engines, intelligent healthcare, and financial risk control (jang2017kbqa; wu2024medical; chen2025ifqa), it has garnered extensive attention in recent years. Although KBQA performs well on simple questions, it still struggles with complex questions in real-world scenarios. This difficulty primarily arises from generalizing to a wide range of unseen complex questions and adapting to the heterogeneous structures of KBs.
@@
Existing semantic-parsing-based (shu2022tiara; jiangunikgqa) and information-retrieval-based KBQA approaches (liu2022graftnet; sun2019pullnet) rely on predefined workflows (e.g., generate-then-retrieve or retrieve-then-rank pipelines), resulting in error propagation and adaptability limitations. Recently, a series of approaches has adopted an agentic reasoning paradigm. As illustrated in the middle of Fig. 1, under this paradigm, LLMs autonomously perform task planning, tool calling, and ultimately derive precise answers through multi-turn interactions with KBs. Specifically, existing agentic approaches fall into two categories, i.e., prompting-based (xiong2024interactive; sunthink) and finetuning-based approaches (jiang2024kgagent; fang2024dara; zhu2025knowagent). The former uses hand-crafted few-shot prompts to guide reasoning, relying on the intrinsic reasoning ability of LLMs. However, few examples are inadequate for the model to understand complex questions, perceiving the KB structure, resulting in low performance. In response, finetuning-based approaches adapt open-source LLMs by constructing agentic reasoning trajectories and fine-tuning the model to follow them. However, most of them rely on process supervision. Since manually creating high-quality trajectories is labor-intensive, they typically synthesize reasoning trajectories by decomposing gold logical forms (e.g., SPARQL). Besides, they learn from trajectories via Supervised Fine-Tuning (SFT) to employ the model with agentic reasoning ability.
@@ 2.1 Traditional KBQA methods
Traditional KBQA methods can be categorized into two primary kinds of approaches: Information Retrieval-based (IR-based) and Semantic Parsing-based (SP-based) ones. IR-based methods (miller2016kvmem; liu2022graftnet; sun2019pullnet; saxena2020embd; shi2021transfernet; zhang2022subgraph) first retrieve query-relevant information from the KB and subsequently rank candidate answers. Among IR-based methods, some (bordes2015large; chen2019bidirectional) incorporate a long-term memory component that can be read from and written to, or perform a path-walking process over the KB to derive the answer (qiu2020stepwise; ren2021lego). SP-based methods (sun2020sparqa; chen2019uhop; lanquery; jiangunikgqa; ye2022rng; shu2022tiara) transform the natural language question into a logical query, link the entities and relations, and execute the completed query to obtain the answer. With the advent of LLMs, they have been employed to enhance both approaches, serving as powerful subgraph retrievers (zhao2024kg; liu2024dual) or semantic parsers (li2023few; nie2024code). However, these enhancements do not fundamentally alter the predefined, multi-stage workflow, which brings inflexibility and error propagation.
@@ 2.2 Agentic KBQA methods
Recently, a new paradigm of agentic reasoning has emerged, which views the LLM as an agent that uses tools to interact with the KB, progressively reason, and finally find the answer. We call them agentic methods in this paper. They can be divided into two main categories: prompt-based and finetuning-based ones. The former methods, such as Interactive-KBQA (xiong2024interactive), QueryAgent (huang2024queryagent), Readi (cheng2024call), and TOG (sunthink), leverage carefully designed interaction strategies and prompts to elicit the model’s reasoning capabilities. However, these methods struggle to equip the model with sufficient agentic capability to effectively perceive the KB and execute complex reasoning. Their success is also heavily dependent on the intrinsic capabilities of LLMs, most of which are closed-source and have an extremely large number of parameters. However, the limited number of examples is insufficient for the model to fully comprehend complex questions and capture the underlying KB structure, leading to suboptimal performance. Finetuning-based methods aim to distill reasoning abilities into smaller, open-source LLMs. Approaches (gu2023don; jiang2024kgagent; zhu2025knowagent; fang2024dara) achieve this by constructing high-quality reasoning trajectories and finetuning on them. However, these trajectories, by decomposing and augmenting gold program annotations and the reliance on process supervision via SFT, limit the model’s capacity for autonomous exploration, resulting in poor robustness and flexibility. Although some recent works have used Monte Carlo Tree Search (MCTS) methods (luokbqao1; xiong2025mcts) to expand samples during inference for incremental fine-tuning, this is an indirect way to foster exploration and still operates within the confines of the process-supervised SFT paradigm. Overall, existing finetuning-based methods fall under SFT-based process supervision, which, as noted in Section 1, struggles to encourage autonomous exploration, leading to less robust and adaptable reasoning trajectories.
@@ 3 Preliminaries
KBQA. Given a natural-language question $q$, the KB $\mathcal{G}$, and a set of topic entities $E_{q}\subseteq\mathcal{E}$ mentioned in $q$, the objective of the KBQA task is to find the answer set that maximizes the conditional probability $P(\mathcal{A}_{q}|q,E_{q},\mathcal{G})$. Notably, following prior work luokbqao1, we assume entity mentions in $q$ are linked to the knowledge graph $\mathcal{G}$ and given as an input.
@@ 4.1 Agentic Initialization
KnowCoder-A1 follows a ReAct-style (yao2023react) agent paradigm that interleaves explicit reasoning and tool use over a KB environment, and terminates with a final answer.
@@
Action Space. At each step, the agent first produces an explicit thinking process and then makes a tool call from a finite action set Action. We instantiate three tools based on Interactive-KBQA (xiong2024interactive):
@@ 4.2 SFT-based Cold-start Stage
Next, to endow the model with stronger exploratory abilities and expand its exploration space, we use a prompt designed to encourage trial-and-error. As the yellow block in Fig. 2 shows, the prompt contains three parts: task description, tool definition, and high-quality demos: a few manually crafted exemplars that demonstrate a general agentic reasoning process, including planning, interacting, and self-correction after errors. Then, we prompt strong models (e.g., Gemini 2.5 Flash  (comanici2025gemini) and DeepSeek-V3  (guo2025deepseek)) to generate multiple candidate trajectories for each question.
@@
  • Exact-match Correctness. The Exact Match (EM) score (talmor2018web) between the predicted and gold answers must be $1$. This ensures that at least one of the predicted answers appears in the gold answers, which largely guarantees the correctness of the trajectory.
@@ 4.3 RL-based Exploration Stage
Based on the cold-start model, in the RL stage, we aim to further promote autonomous exploration and eliminate the reliance on process supervision. To achieve this, we employ the outcome-based Group Relative Policy Optimization (GRPO) method (shao2024deepseekmath), where any trajectory leading to a correct answer receives a positive incentive, thereby encouraging the agent to explore and acquire stronger reasoning capabilities. Moreover, to ensure the RL training is both balanced and effective, we implement two key strategies.
@@ 4.3.1 Policy Optimization via GRPO
Inspired by the recent success of policy gradient methods on reasoning tasks (jin2025search; xue2025simpletir; luo2025graph), we employ the GRPO method for optimization. GRPO estimates the baseline using the rewards from a group of rollouts for the same query to compute the advantage. Its objective function can be expressed as follows:
@@
In this paper, we adopt an on-policy variant of GRPO. This variant applies a single SGD step to the entire rollout batch, which sets the importance ratio in the aforementioned formula to 1. This approach has also been shown to be an effective way to incentivize exploration (he2025skywork).
@@ 5.1 Experiment Setup
Dataset and Metrics. We evaluate KnowCoder-A1 on three widely-used KBQA datasets: WebQSP, CWQ, and the generalization-focused GrailQA, using a subset of the official testing data. Following existing work (xiong2024interactive), we primarily evaluate the F1 metrics, supplemented by Random Hits@1(RHits@1) and EM score. As our approach does not rely on semantic parsing, we do not report logical form accuracy (gu2021beyond).
Baselines. We compare KnowCoder-A1 against two main categories of methods: prompting-based and fine-tuning-based methods. For the former, we include a diverse range of prompting strategies, from direct prompting, such as IO, CoT, etc., and agent-based reasoning method Interactive-KBQA (xiong2024interactive). To ensure a fair comparison and reproducibility, we report the scores for these methods directly from their original publications. For the fine-tuning-based methods, we select SOTA methods finetuned on full data, include RnG-KBQA (ye2022rng) and TIARA (shu2022tiara), and the directly comparable low-source methods, which are finetuned on low-source data, includes agentic approaches like SFT-traj. (xiong2024interactive), KBQA-o1 (luokbqao1), and MCTS-KBQA (xiong2025mcts).
@@ A.1 Datasets
We evaluate the reasoning performance on three widely-used datasets: WebQSP (talmor2018web), CWQ (talmor2018web), and GrailQA (gu2021beyond). These datasets are constructed based on Freebase and consist of natural language questions paired with their corresponding SPARQL queries. Specifically, WebQSP features questions where the topic entity and answer entities are at most 2 relation hops apart. CWQ extends WebQSP by incorporating more complex entities and constraints, with relation paths up to 4 hops. In contrast to WebQSP and CWQ, GrailQA is designed for a more comprehensive evaluation of generalization capabilities across three settings: independently and identically distributed (I.I.D.), compositional, and zero-shot.
To ensure consistent results and reduce evaluation overhead, we follow the setting of previous work xiong2024interactive, which uniformly sample questions across different types in WebQSP and CWQ. For the GrailQA dataset, we randomly sample 600 instances for evaluation.
@@ A.2 Baselines
Fine-tuning-based Methods. We consider two settings: fine-tuning on partial data (low-source setting) and on full data. In the low-source setting, we select the agentic approaches, includes the SFT model in Interactive-KBQA (xiong2024interactive), which uses the same toolset as our method (abbreviated as SFT-traj.), two MCTS-based agentic methods, KBQA-o1 (luokbqao1) and MCTS-KBQA (xiong2025mcts). Notably, since our method is trained on only a subset of the data and uses no SPARQL annotations, its performance is directly comparable to this category of baselines. For the full-data setting, which serves as a secondary point of reference, we include the SOTA SP-based methods RnG-KBQA (ye2022rng) and TIARA (shu2022tiara).
Prompting-based Methods. We selected a diverse set of prompting strategies for comparison. These include methods based on direct prompting, Chain-of-Thought (CoT), and self-consistency with CoT; approaches that use an LLM to enhance semantic parsing (KB-BINDER(li2023few) and KB-Coder (nie2024code)) and subgraph retrieval (tian2024augmenting); and Interactive-KBQA (xiong2024interactive) that prompts an LLM to reason as an agent. For a fair comparison, we report the scores for these methods directly from their official papers, as the specific API versions used in their experiments are no longer accessible.
@@ A.3 Evaluation Metrics
We primarily report the F1 score since the answers are returned as an unordered list. Following previous work (xiong2024interactive), we also report Random Hits@1 (RHits@1) and Exact Match (EM) for reference. As our method is not fully based on semantic parsing, we do not report the semantically equivalent of logical forms (another score commonly abbreviated as EM) in  gu2021beyond.